a cobbler makes slippers for bats and cats
today he makes twice as many bats as cats
that is 104 individual slippers
how many bat slippers did he make

Answer by macston(5194)

C=Number cat slippers B= Number bat slippers=2C
C+B=C+2C=104 (total slippers)
3C=104
C=104/3= 34 2/3 He made 34 2/3 Cat slippers
B=2C=2(34 2/3)=69 1/3
ANSWER He made 69 1/3 Bat slippers
CHECK
Bat slippers + cat slippers = total slippers
69 1/3+ 34 2/3=104
104=104